OpenStudy (anonymous):

1. 0.817643 = rational or irrational 2. -√88 rational or irrational 3. -√169 rational or irrational 4. Estimate √58 to the nearest integer. Help! Thank you!

OpenStudy (hba):

Do you have an intuition about rational and irrational number ?

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

What is that supposed to mean

OpenStudy (hba):

What is your mother toungue ?

OpenStudy (anonymous):

That makes no sense

OpenStudy (hba):

Let me rephrase it for you, Do you know what a rational and irrational number is ?

OpenStudy (aravindg):

any number with terminating decimals is a rational number ,but those numbers having not-terminating(maybe repeating) decimals are irrational numbers.

OpenStudy (anonymous):

Yes I do

OpenStudy (hba):

Is pi a rational or irrational number ?

OpenStudy (anonymous):

irrational

OpenStudy (hba):

What about 0.5 ?

OpenStudy (anonymous):

rationsl

OpenStudy (aravindg):

good :)

OpenStudy (aravindg):

i hope you understand it now :)

OpenStudy (hba):

So, 0.817643 is rational or irrational ?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

rational

OpenStudy (hba):

See, 0.5 =1/2 so rational.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Thanks anyways I got it

OpenStudy (phi):

rational means you can write it down as a decimal. irrational means the decimals keep going, never repeating, no rhyme or reason for what the next decimal will be pi is a good example. roots (in general) are irrational unless they are roots of perfect squares. example square root of 4 is 2 but sqrt(5) is irrational. btw, same idea for cube roots, etc...

OpenStudy (anonymous):

Thankies!! :)

OpenStudy (phi):

The trick part in these questions is recognizing that a number is a perfect square. 169 is 13*13, so the square root of 169 is 13, and rational

OpenStudy (phi):

88 is 2*44 or 2*2*2*11 and so sqrt(88)= sqrt(2*2*2*11)= sqrt(2*2)*sqrt(2*11)= 2sqrt(22) and sqrt(22) is irrational, and 2 times an irrational is still irrational.

OpenStudy (phi):

4. Estimate √58 to the nearest integer. Are you allowed a calculator? Or did they give you a way to do this problem?

OpenStudy (anonymous):

calculator

OpenStudy (phi):

find the square root, and round to the nearest whole number. to round, look at the digit in the tenths place (next to the decimal point) if 4 or less, drop the decimals and that is the answer if 5 or bigger, drop the decimals, but ADD 1 to the answer example: 100.3 to the nearest integer is 100 (3 is 4 or less, so drop the decimals) 210.50013 to the nearest integer is 210+1 (because 5 is 5 or bigger) you get 211

OpenStudy (anonymous):

Alrighty thank you

OpenStudy (phi):

if you are not allowed a calculator, write down the squares near 58 7*7 is 49 58 is between them 8*8 is 64 58 is 9 away from 49 and 6 away from 64. so 58 is closer to 64 than it is to 49 that is a clue that sqrt(58) is closer to 8 than to 7